Subtract. Write the answer in lowest terms.

17/24 - 1/24

Answer by CharlesG2(834)

Subtract. Write the answer in lowest terms.

17/24 - 1/24

the denominators (the lower parts of the fractions are the same)
17 - 1 = 16

16/24 = 2/3 = answer in lowest terms
(divided top and bottom by 8)